# Artificial Inductive Bias for Synthetic Tabular Data Generation in Data-Scarce Scenarios

### v1

@@ 3.4 Results
Table 2: News dataset JS and KL results for each Scenario. Big data represents the ideal case where many samples ($N=10,000$) are available to generate reliable synthetic data. Low data represents a more realistic scenario in which a limited number of samples ($N=300$) are available, posing a challenge for synthetic data generation. The next rows compare the divergences obtained by each methodology (pre-training, model averaging, MAML, and DRS) applied to the low data scenario. Values in bold indicate an improvement due to the technique used. Results are represented as mean (std). Lower is better.
Table 3 further reinforces the efficacy of the methodologies proposed in the King dataset. The VAE model consistently improves on the lower bounds established for the JS and KL divergences across all techniques. For CTGAN, the results for the JS and KL divergences are consistent with the lower bounds. This suggests that while CTGAN does not produce significant reductions in divergence metrics, it appears to maintain the quality of the data distribution compared to the lower bounds. However, it should be noted that CTGAN exhibits consistently lower gains across datasets compared to the VAE. This may be attributed to inherent GAN framework instabilities.
Scenario N VAE JS CTGAN JS VAE KL CTGAN KL
Big data 10000 0.862 (0.002) 0.777 (0.003) 4.768 (0.072) 3.124 (0.115)
Low data 300 0.927 (0.002) 0.940 (0.003) 13.763 (0.696) 7.470 (0.392)
[…]
AVG 300 0.740 (0.002) N/A 3.489 (0.209) N/A
MAML 300 0.910 (0.002) N/A 6.436 (0.496) N/A
DRS 300 0.809 (0.003) N/A 4.321 (0.215) N/A
Table 3: King dataset JS and KL results for each Scenario. Big data represents the ideal case where many samples ($N=10,000$) are available to generate reliable synthetic data. Low data represents a more realistic scenario in which a limited number of samples ($N=300$) are available, posing a challenge for synthetic data generation. The next rows compare the divergences obtained by each methodology (pre-training, model averaging, MAML, and DRS) applied to the low data scenario. Values in bold indicate an improvement due to the technique used. Results are represented as mean (std). Lower is better.
The results obtained using the Intrusion dataset are summarized in Table 4. These values align with the findings for the News dataset, demonstrating consistent improvements in divergence metrics for reliable case ($M=7500$ and $N=1000$) across different methodologies. Similarly to the News dataset, Intrusion presents a high number of features, resulting in higher dimensionality. This increased dimensionality poses a greater challenge in generating synthetic data that closely resembles the real-world data distribution. Consequently, the divergence results for Intrusion either maintain or exhibit smaller improvements compared to lower-dimensional datasets. Despite the challenges posed by the high dimensionality of Intrusion, the proposed methodologies still demonstrate their effectiveness in improving divergence metrics, particularly for reliable cases.
Scenario N VAE JS CTGAN JS VAE KL CTGAN KL
Big data 10000 0.760 (0.013) 0.531 (0.033) 2.744 (0.084) 2.623 (0.537)
Low data 300 0.920 (0.003) 0.961 (0.002) 6.216 (0.154) 8.841 (0.710)
[…]
AVG 300 0.867 (0.007) N/A 5.798 (0.295) N/A
MAML 300 0.913 (0.003) N/A 6.359 (0.054) N/A
DRS 300 0.835 (0.009) N/A 4.587 (0.166) N/A
Table 4: Intrusion dataset JS and KL results for each Scenario. Big data represents the ideal case where many samples ($N=10,000$) are available to generate reliable synthetic data. Low data represents a more realistic scenario in which a limited number of samples ($N=300$) are available, posing a challenge for synthetic data generation. The next rows compare the divergences obtained by each methodology (pre-training, model averaging, MAML, and DRS) applied to the low data scenario. Values in bold indicate an improvement due to the technique used. Results are represented as mean (std). Lower is better.
Finally, Table 5 summarizes the results of applying various methodologies to generate synthetic tabular data using the VAE model. Across different methodologies, the results consistently show positive gains, except for MAML. These gains are computed as the difference between the metric in the lower bound and any of the proposed methods. This indicates that the proposed techniques generate synthetic data that are closer to the real-world data distribution than in the case of not using them. This improvement is particularly evident for the model averaging methodology, which consistently outperforms other techniques in terms of divergence reduction. MAML’s performance might be limited in this scenario due to its reliance on a large number of seeds. Since the results presented do not specify the number of seeds used, MAML might not have had enough to learn effectively [49]. The consistent improvement in divergence metrics highlights the robustness and generalizability of the proposed techniques. These findings suggest that our approach is an effective methodology for generating high-quality synthetic tabular data that can be used for various applications.
@@
Adult 0.159 (0.482) 0.271 (0.388) 0.173 (0.525) 0.317 (0.455) 0.030 (0.092) 0.011 (0.016) 0.142 (0.430) 0.271 (0.388)
News 0.093 (0.111) 1.065 (0.233) 0.230 (0.274) 1.985 (0.433) -0.011 (-0.013) -0.594 (-0.130) 0.194 (0.231) 2.133 (0.465)
King 0.064 (0.070) 8.477 (0.616) 0.187 (0.202) 10.274 (0.746) 0.017 (0.018) 7.327 (0.532) 0.118 (0.128) 9.442 (0.686)
Intrusion 0.127 (0.138) 2.385 (0.384) 0.053 (0.057) 0.419 (0.067) 0.006 (0.007) -0.143 (-0.023) 0.085 (0.092) 1.629 (0.262)
[…]
Table 5: Gains using the proposed methodology for the VAE. Gains are represented in the following format: absolute gain (relative gain). The methodology achieves relative gains of up to 50% in JS divergence, which is bounded, and up to 75% in KS divergence. Bold values indicate positive gain. Higher is better.
It is important to acknowledge the varying computational demands of the compared methods. Model averaging emerges as the most efficient approach, as inductive bias generation only involves calculating a mean, resulting in minimal computational overhead. In contrast, MAML exhibits the highest computational load due to its intricate optimization procedure. Pre-training and DRS fall between these two extremes, both requiring the training of a DGM to establish the inductive bias. Considering these findings alongside the results presented in Table 5, we recommend against using MAML. It offers minimal performance gains with significant computational cost. The other methods, on the other hand, provide a more favorable trade-off between computational efficiency and performance. Furthermore, as detailed in the Appendix, reliably quantifying the benefits of our methodology in a realistic, limited-data setting is challenging. This implies that validation with a large number of samples is necessary to definitively assess which of our proposed inductive bias techniques yields superior results for a specific dataset. However, the experimental results strongly suggest potential gains that warrant further exploration.
@@ 4 Conclusions
This research proposed a novel approach to generate synthetic tabular data using DGMs in the context of limited datasets. Our approach leverages four distinct techniques to artificially introduce an inductive bias that guides the DGM towards generating more realistic and informative synthetic data samples. These techniques encompass two transfer learning approaches: pre-training and model averaging; and two meta-learning approaches: MAML and DRS. To facilitate the application of model averaging, MAML, and DRS, we employ the VAE model from [15] and train multiple instances with different random seeds. This allows us to leverage the ensemble properties of the VAE models for techniques like model averaging and further enables the application of meta-learning algorithms like MAML and DRS. We also used the CTGAN [7] to assess pre-training so that we can compare other architectures of well-known models for synthetic tabular data generation. We used divergence metrics, in particular JS and KL divergences, to compare the real and synthetic data distributions generated. The experimental results consistently demonstrate the effectiveness of our proposed approach in generating high-quality synthetic tabular data, particularly when using transfer learning techniques. These techniques significantly improve the divergence metrics, indicating a closer resemblance between the synthetic and real data distributions. Our approach offers several advantages over existing methods. Firstly, it effectively addresses the challenge of generating realistic synthetic data from small datasets, a common limitation in many real-world applications. Secondly, the use of transfer learning and meta-learning techniques enhances the inductive bias of the DGM, leading to more meaningful and informative synthetic data samples. However, it is also important to acknowledge the trade-offs associated with our methodology. Training VAEs with these techniques requires training multiple VAE models with different random seeds. This can lead to a significant increase in computational cost compared to simpler DGM training methods. While divergence metrics provide a valuable measure of distributional similarity, their ability to reliably assess the improvement in synthetic data quality for specific downstream tasks can be limited, especially with small datasets, as detailed in the Appendix. Nonetheless, our methodology may provide significant gains in JS divergence, of up to 50% according to our experimental results.
In conclusion, our proposed approach provides a promising solution for generating high-quality synthetic tabular data from small datasets, particularly when VAEs are employed to apply transfer learning techniques. We believe that this work has the potential to make a significant contribution to the field of synthetic data generation and machine learning applications that rely on small datasets. However, there are several research lines to be addressed. While the current study focuses on VAEs and GANs, investigating the applicability of our framework to other DGM architectures could provide valuable insights. In addition, our current approach does not explicitly incorporate domain knowledge. Future research could explore mechanisms to integrate domain-specific information from an expert into the inductive bias generation process, potentially leading to even more realistic and informative synthetic data. Lastly, although divergence metrics offer a valuable measure of distributional similarity, exploring additional evaluation techniques that assess the quality and usefulness of synthetic data for specific downstream tasks would provide a more comprehensive understanding of the effectiveness of our methodology, valid also for the case when little amount of data is available for validation, which is a current limitation of this work.
@@ Acknowledgements
Appendix A Appendix

### v2

@@ 3.4 Results
Table 1: JS and KL results for each scenario. ‘Big data’ represents the ideal case with ample samples ($N=10,000$) for reliable synthetic data generation. In contrast, ‘Low data’ simulates a more constrained setting ($N=300$). The next rows compare divergences and utility validation metrics obtained with different methodologies (pre-training, model averaging, MAML, and DRS) in the ‘Low data’ setting. Results are shown as mean (std). For divergence metrics (JS, KL), lower values indicate better performance. For utility metrics, the interpretation depends on the task: for classification tasks (accuracy), higher values are better; for regression tasks (MSE), lower values are better. Bold values denote improvements due to the technique.
Table˜2 reports the gains obtained when applying the proposed methodologies to the VAE model, measured as the absolute and relative improvement in JS and KL divergence compared to the baseline ‘Low data’ scenario. The results show consistent improvements across most methods and datasets, with the most significant and robust gains achieved using the ‘AVG’ strategy. While the ‘Pre-train’ and ‘DRS’ techniques also show positive and consistent gains in both divergence metrics, the performance of ‘MAML’ is notably weaker and less stable. Although ‘MAML’ yields some improvement in the Adult dataset, it fails to generalize across others and, in some cases, degrades KL divergence (e.g., King). This limited performance may stem from the nature of ‘MAML,’ which requires a diverse set of tasks for effective meta-learning—something not fully met with only 10 training seeds per dataset. Overall, these findings confirm that the proposed inductive bias strategies enhance the fidelity of synthetic data generation, with ‘AVG,’ ‘Pre-train,’ and ‘DRS’ emerging as the most effective techniques.
@@
Adult 0.238 (0.567) 0.601 (0.593) 0.257 (0.612) 0.624 (0.615) 0.062 (0.149) 0.138 (0.136) 0.240 (0.572) 0.551 (0.544)
King 0.043 (0.046) -0.252 (-0.042) 0.195 (0.212) 2.641 (0.443) 0.020 (0.022) -4.949 (-0.830) 0.048 (0.052) -0.155 (-0.026)
Letter 0.069 (0.106) 0.294 (0.148) 0.073 (0.112) 0.243 (0.122) -0.015 (-0.023) -0.013 (-0.007) 0.046 (0.070) 0.352 (0.177)
Average 0.117 (0.227) 0.214 (0.233) 0.175 (0.312) 1.169 (0.393) 0.023 (0.049) -1.608 (-0.234) 0.111 (0.231) 0.249 (0.232)
Table 2: Gains using the proposed methodology for the VAE. Gains are represented in the following format: absolute gain (relative gain). The methodology achieves relative gains of up to 60% in JS and KL divergences. Bold values indicate positive gain. Higher is better.
Interestingly, the accuracy achieved when training with a limited number of real samples closely approaches that obtained under the ‘Big data’ configuration. This suggests that a small subset of the data may retain the most critical discriminative information required for the task. Furthermore, even in scenarios where divergence metrics indicate substantial discrepancies in the joint distribution, the utility metric (measured via accuracy) remains high. This observation implies that the generative model is effectively capturing and reproducing key informative features, which are sufficient for maintaining classification performance despite broader distributional misalignments.
Finally, it is important to acknowledge the varying computational demands of the compared methods. Model averaging is the most efficient approach, as inductive bias generation only involves calculating a mean, resulting in minimal computational overhead. In contrast, ‘MAML’ exhibits the highest computational load due to its intricate optimization procedure. ‘Pre-train’ and ‘DRS’ fall between these two extremes, both requiring the training of a DGM to establish the inductive bias. Considering these findings alongside the results presented in Table˜2, we recommend against using ‘MAML.’ It offers minimal performance gains with significant computational costs. The other methods, on the other hand, provide a more favorable trade-off between computational efficiency and performance. Furthermore, as detailed in A, reliably quantifying the benefits of our methodology in a realistic, limited-data setting is challenging. This implies that validation with a large number of samples is necessary to definitively assess which of our proposed inductive bias techniques yields superior results for a specific dataset. However, the experimental results strongly suggest potential gains that warrant further exploration.
@@ 4 Conclusions
This research proposed a novel approach to generate synthetic tabular data using DGMs in the context of limited datasets. Our approach leverages four distinct techniques to artificially introduce an inductive bias that guides the DGM toward generating more realistic and informative synthetic data samples. These techniques encompass two transfer learning approaches — pre-training and model averaging — and two meta-learning approaches: MAML and DRS. To facilitate the application of model averaging, MAML, and DRS, we employ the VAE model from apellaniz2024improved and train multiple instances with different random seeds. This enables us to leverage the ensemble properties of the VAE models for techniques such as model averaging and further facilitates the application of meta-learning algorithms like MAML and DRS. We also utilized the CTGAN (xu2019modeling, ) to evaluate pre-training and compare it with other well-known models’ architectures for generating synthetic tabular data. Our approach offers several advantages over existing methods. Firstly, it effectively addresses the challenge of generating realistic synthetic data from small datasets, a common limitation in many real-world applications. Secondly, the use of transfer learning and meta-learning techniques enhances the inductive bias of the DGM, leading to more meaningful and informative synthetic data samples. However, it is also important to acknowledge the trade-offs associated with our methodology. Training VAEs with these techniques requires training multiple VAE models with different random seeds. This can lead to a significant increase in computational cost compared to simpler DGM training methods. While divergence metrics provide a valuable measure of distributional similarity, their ability to reliably assess the improvement in synthetic data quality for specific downstream tasks can be limited, especially with small datasets, as detailed in A.
In conclusion, our approach offers a promising solution for generating high-quality synthetic tabular data from small datasets, particularly when VAEs apply transfer learning techniques. This work contributes to the generation of synthetic data and machine-learning applications that rely on limited data. Building on this foundation, future research should extend our methodology to other DGMs, such as diffusion models or normalizing flows, which may require adapting the bias injection techniques to new architectures. It would also be beneficial to integrate domain-specific priors or structural knowledge, especially in fields such as medicine or engineering, where expert information can enhance the fidelity of the generated results. Improving evaluation strategies beyond distributional metrics (particularly under low validation budgets) remains a key challenge, and incorporating task-aware or uncertainty-aware metrics could be a promising direction. Lastly, automating hyperparameter tuning and seed selection through meta-configuration tools or Bayesian optimization could reduce manual intervention and improve reproducibility. Overall, our work presents a general and extensible framework for enhancing synthetic data generation in low-resource tabular settings, offering a practical solution for real-world applications that require privacy-preserving and high-fidelity data synthesis.
Altogether, our methodology offers a flexible and extensible solution for generating high-fidelity synthetic data under small-sample constraints. By combining principled inductive bias strategies with scalable generative modeling, we lay the groundwork for future developments in privacy-preserving, data-efficient machine learning across domains where real data are scarce or sensitive.
@@ Acknowledgements
Appendix A Additional Experiments under Constrained Validation Settings